# takeawaybox backgroundcolor=gray!30, shadow=false, innerleftmargin=5pt, innerrightmargin=5pt, roundcorner=10pt, skipabove=12pt

### v3

@@ 12.2.2 Challenges in Representation Operationalization
Assumptions About Models’ Representations. Methods for identifying concept operators come with assumptions about the shape and geometry of representations in the models’ activations. Common assumptions about representations are that concepts are represented as linear direction (Qiu et al., 2024), that representations do not change throughout a sequence of inputs, that they are the same across context, that they are localized in a single layer, and that they do not rely on interactions between layers. However, if the true form of the representation does not match these assumptions, it will lead to unsuccessful or less effective Representation Identification and Control.
These assumptions have been criticized (Luo et al., 2024; Chu et al., 2024). Assumptions have been changed and weakened by shifting from a point-in-space to a direction-magnitude view (Pham & Nguyen, 2024), adapting the strength and direction of steering to the context (Wang et al., 2024c), or applying non-linear operations (Qiu et al., 2024; Hoscilowicz et al., 2024).
@@ 12.2.3 Challenges in Representation Control
{mdframed}
[style=takeawaybox] Takeaway: Challenges in Representation Engineering RepE struggles with steering multiple concepts at once, controlling long-form generations, and providing a reliable steering effect. These weaknesses arise because of challenges like the inability to disentangle spuriously correlated concepts, to posit correct assumptions about models’ representations, and to retain the models’ natural distribution of activations.
@@ 13.5 Building a More Rigorous Science of Representation Engineering
The field of Representation Engineering is young and does not yet stand on a strong scientific foundation. To enable progress in the field, we point out the opportunity to make research into RepE more rigorous and to provide a better grounding for the field.
13.5.1 Best Practices for Evaluating RepE Methods
Better Measurements of Controllability. It is, in general, hard to evaluate whether the outputs of an LLM were successfully steered towards a concept. Many papers, therefore, employ LLMs to judge how much an output corresponds to a concept. While this can be a convenient and effective metric (Zheng et al., 2023), using LLMs-as-a-judge is known to be biased and sometimes inaccurate (Fu et al., 2024; Zheng et al., 2023). Alternatively, many studies rely on existing benchmarks, which largely only include short answers, thus ignoring the evaluation of more free- and long-form generation.
Reporting Changes in Models’ Quality. Only a minority ( 35%) of papers report the change in models’ capabilities after applying Representation Engineering. Since we know that RepE can reduce these capabilities, it is important to have a standard of reporting measures of LLMs’ quality so that we can compare the merits of each method. These could include performance on popular performance benchmarks or changes in perplexity Braun et al. (2024).
Measuring Out-of-Distribution Generalisation. As noted earlier, measuring how effectively an operator is able to control an LLM in cases that are different from those where the operator was identified is an open challenge. However, most papers only measure performance on distributions that are similar to the training data, which leads to overly optimistic results.
Comparing to SOTA Methods. When proposing a new method, its effectiveness should be compared to the current state-of-the-art method. However, often papers only compare to early methods that have since been surpassed.
Testing on Different Model Sizes. Appendix A.1 finds that most RepE experiments are conducted on models between 3-10 billion parameters. To prove the general effectiveness of a RepE method, it is necessary to evaluate it on models of different sizes. Specifically, larger models need to be tested to prove that the method scales to frontier models.
13.5.2 Benchmarking RepE Pipelines
There is currently no benchmark that provides a way to comprehensively evaluate the effectiveness of a RepE pipeline. Thus we are currently unsure which pipeline is most effective in which situations. A benchmark can also serve as a guide for the future development of RepE methods by providing clarity about the effectiveness of methods and being a target to aim for. Such unified evaluations can also surface previously undiscovered limitations of methods (Brumley et al., 2024).
Such a benchmark should cover a range of concepts of different types since different RepE methods might steer different concepts more effectively (Brumley et al., 2024). Hereby, the focus should lie on use cases where RepE is foreseen to be practically useful, such as improving truthfulness, changing goals, adapting situational awareness, or removing societal biases. Furthermore, evaluations should cover multiple models, test the influence of dataset size, and show the impact of different steering strengths. Additionally, it should specifically evaluate the weaknesses of current RepE methods (see Section 12.1), thus putting a spotlight on improvements on these challenges. Of course, this benchmark should make sure to follow best practices outlined above (Section 13.5.1.
AxBench (Wu et al., 2025) is a benchmark that evaluates 9 different RepE methods on 500 concepts. However, all concepts are of a similar type, testing whether a concrete, narrow topic is contained in a text. Extensions to this benchmark should consider different types of concepts. Im & Li (2025) conduct an evaluation of different Input Reading methods for Representation Identification. However, they only focus on high-level behaviors and always use a vector as a concept operator and Linear Addition as a steering function. An extension could consider different concepts and methods for Representation Control.
13.5.3 Groundwork for RepE
@@
{mdframed}
[style=takeawaybox] Takeaway: Opportunities for Future Research Future RepE methods can expand restrictive notions of representations, combine multiple RI methods and attempt more complex interventions. Untapped applications lie in steering goals and values of agents or preventing deceptive behavior. To progress the field of RepE, thorough evaluations and best practices are needed.
@@ A.1 Commonly Used Models
[…]
Figure 4 shows that a majority of models to which RepE has been applied on lie between $3-10$ billion parameters. This is underlined by the fact that 9 out of the 10 most popular models for RepE are all between $7-13$ billion parameters. Furthermore, 8 of them belong to the Llama family of models or are derived from Llama models. This calls for more diversity in which model experiments are performed. However, the model scale might not be a crucial factor for the effectiveness of RepE. Across 4 papers that compare the steerability of models across different numbers of parameters within the same model family, none find a clear relationship between scale and steerability (Arditi et al., 2024; Xu et al., 2024a; Arora et al., 2024; Bortoletto et al., 2024). Furthermore, only Templeton et al. (2024) are able to experiment on a proprietary, state-of-the-art model, thus leaving open questions about how effective RepE is at larger scales and for the most capable models.
@@ A.3 Statistics about Publications on Representation Engineering
{mdframed}
[style=takeawaybox,nobreak=true] Takeaway: Meta-survey
  • Models: A majority of models used have between 3-10 billion parameters. According to current literature, there appears to be no relationship between RepE effectiveness and model scale.
  • Samples: More samples make RepE more effective.
  • Publication statistics: A majority of papers on RepE come from academics. The literature is very new but growing rapidly.
@@ Appendix D Papers that compare RepE to Prompting, Fine-tuning and Decoding-based methods
Prompting. Zhang et al. (2024c); Cai et al. (2024); Konen et al. (2024); Arditi et al. (2024); Ghandeharioun et al. (2024); Zhang & Viteri (2024); Chen et al. (2024b); Stickland et al. (2024); Guo et al. (2024); Qiu et al. (2024); Leong et al. (2023); Li et al. (2024c); Luo et al. (2024); Cao et al. (2024b); Scalena et al. (2024); Koo et al. (2024); Wang et al. (2024b); Liu et al. (2024b); Li et al. (2024e); Liu et al. (2024a); Wang et al. (2024c); Li et al. (2024a)
Fine-tuning. Lucchetti & Guha (2024); Li et al. (2024d); Yang et al. (2024); Price et al. (2024); Pham & Nguyen (2024); Liu et al. (2024b); Wang et al. (2024b); Koo et al. (2024); Wu et al. (2024b); Yu et al. (2024); Leong et al. (2023); Qiu et al. (2024); Guo et al. (2024); Stickland et al. (2024); Qian et al. (2024); Chen et al. (2024b); Zhang et al. (2024b); Liu et al. (2023); Wang et al. (2024a); Ackerman (2024); Li et al. (2024b); Yin et al. (2024); Cao et al. (2024a); Wu et al. (2024a; c)

### v4

@@ 13.2.2 Challenges in Representation Operationalization
Assumptions About Models’ Representations. Methods for identifying concept operators come with assumptions about the shape and geometry of representations in the models’ activations. Common assumptions about representations are that concepts are represented as linear direction (Qiu et al., 2024), that representations do not change throughout a sequence of inputs, that they are the same across context, that they are localized in a single layer, and that they do not rely on interactions between layers.
However, if the true form of the representation does not match these assumptions, it will lead to unsuccessful or less effective Representation Identification and Control. If a concept representation is non-linear, a linear concept operator can at most be a first-order approximation of the true representation and thus provides less precise and effective steering. Similarly, when the true representation of a concept depends on the context or can change throughout trajectories, steering with one static concept operator is at best optimal in a subset of situations. Additionally, when concepts rely on interactions between layers, single layer steering cannot capture all aspects of that concept or could fail to correctly influence the concept in later layers.
These assumptions have been criticized (Luo et al., 2024; Chu et al., 2024). See also our discussion of the validity of the Linear Representation Hypothesis in Section 5.1. Assumptions have been changed and weakened by shifting from a point-in-space to a direction-magnitude view (Pham & Nguyen, 2024), adapting the strength and direction of steering to the context (Wang et al., 2024c), or applying non-linear operations (Qiu et al., 2024; Hoscilowicz et al., 2024).
@@ 13.2.3 Challenges in Representation Control
Takeaway: Challenges in Representation Engineering
RepE struggles with steering multiple concepts at once, controlling long-form generations, and providing a reliable steering effect. These weaknesses arise because of challenges like the inability to disentangle spuriously correlated concepts, to posit correct assumptions about models’ representations, and to retain the models’ natural distribution of activations.
@@ 14.5 Building a More Rigorous Science of Representation Engineering
The field of Representation Engineering is young and does not yet stand on a strong scientific foundation. To enable progress in the field, we point out the opportunity to make research into RepE more rigorous and to provide a better grounding for the field. The most important development here would be the introduction of a high-quality, widely accepted benchmark for measuring RepE methods, which we propose in Section 8.3 as well as the adoption of rigorous evaluation standards discussed in Section 8.2.
@@
Takeaway: Opportunities for Future Research
Future RepE methods can expand restrictive notions of representations, combine multiple RI methods and attempt more complex interventions. Untapped applications lie in steering goals and values of agents or preventing deceptive behavior. To progress the field of RepE, thorough evaluations and best practices are needed.
@@ 15 Conclusion
15.1 Limitations
Our survey is only comprehensive with regard to RepE papers published before 31.8.2024. While we include many papers published until 31.2.2025 we do not comprehensively cover them all. Papers covered after this period are not covered.
While we attempted to synthesis the quantitative evidence on best practices for RepE in Section 7.2, we had to leave many questions unanswered. This is due to a lack of empirical comparisons that shed light on these questions.
Broader Impacts
RepE is a general tool for controlling the behavior of LLMs. In many cases RepE has been devised and used for applications such as steering LLMs to reduce harmful outputs, align them with human intent, make them more honest or uncover societal biases. However, RepE faces a dual-use risk since these concepts can also be steered negatively. For example RepE has been used to jailbreak LLMs to bypass safety mechanisms and produce more harmful outputs. Similarly, RepE could be used to steer the model to become misaligned, generate deceptive text or to strengthen societal biases.
Additionally, RepE has been used to interpret model internals. However, it does not offer perfect explainability. Thus there is a risk that an illusion of interpretability is created where users trust the these explanations more than is warranted.
@@ A.1 Commonly Used Models
We find that a majority of experiments ($164\text{ out of }229$) are performed on base models that have not been trained with RLHF, while $65$ experiments are conducted on preference-tuned models. While experimental results in Dong et al. (2024b); von Rütte et al. (2024) find a larger effect from applying RepE to base models than to instruction-tuned models, Bortoletto et al. (2024); Wang et al. (2024c) find no clear relationship between instruction-tuning and steering effectiveness.
Figure 4 shows that a majority of models to which RepE has been applied on lie between 3103-10 billion parameters. This is underlined by the fact that 9 out of the 10 most popular models for RepE are all between 7137-13 billion parameters. Furthermore, 8 of them belong to the Llama family of models or are derived from Llama models. This calls for more diversity in which model experiments are performed. However, the model scale might not be a crucial factor for the effectiveness of RepE. Across 4 papers that compare the steerability of models across different numbers of parameters within the same model family, none find a clear relationship between scale and steerability (Arditi et al., 2024; Xu et al., 2024a; Arora et al., 2024; Bortoletto et al., 2024). Furthermore, only Templeton et al. (2024) are able to experiment on a proprietary, state-of-the-art model, thus leaving open questions about how effective RepE is at larger scales and for the most capable models.
@@ A.3 Statistics about Publications on Representation Engineering
Takeaway: Meta-survey Models: A majority of models used have between 3-10 billion parameters. According to current literature, there appears to be no relationship between RepE effectiveness and model scale. Samples: More samples make RepE more effective. Publication statistics: A majority of papers on RepE come from academics. The literature is very new but growing rapidly.
@@ Appendix D Papers that compare RepE to Prompting, Fine-tuning and Decoding-based methods
Prompting. Zhang et al. (2024c); Cai et al. (2024); Konen et al. (2024); Arditi et al. (2024); Ghandeharioun et al. (2024); Zhang & Viteri (2024); Chen et al. (2024b); Stickland et al. (2024); Guo et al. (2024); Qiu et al. (2024); Leong et al. (2023); Li et al. (2024c); Luo et al. (2024); Cao et al. (2024b); Scalena et al. (2024); Wang et al. (2024b); Liu et al. (2024b); Li et al. (2024e); Liu et al. (2024a); Wang et al. (2024c); Li et al. (2024a)
Fine-tuning. Lucchetti & Guha (2024); Li et al. (2024d); Yang et al. (2024); Price et al. (2024); Pham & Nguyen (2024); Liu et al. (2024b); Wang et al. (2024b); Wu et al. (2024b); Yu et al. (2024); Leong et al. (2023); Qiu et al. (2024); Guo et al. (2024); Stickland et al. (2024); Qian et al. (2024); Chen et al. (2024b); Zhang et al. (2024b); Liu et al. (2023); Wang et al. (2024a); Ackerman (2024); Li et al. (2024b); Yin et al. (2024); Cao et al. (2024a); Wu et al. (2024a; c)
@@
D.1 Table of empirical comparisons
Furthermore, we list the empirical comparisons between RepE and prompting and fine-tuning methods in Table 14 and Table 15 respectively. Here we include the steering effect and, if available, a measure for model quality for each concept that was compared. When scores for multiple models or subsets of the test dataset are given we average them into one number. Lastly, these table do not include all papers listed above, since we only select papers that report numbers and leave out papers that solely show performance on a graph. This is because reading values of a graph can be unprecise.
Table 14: Collection of studies that compare the steering effect and model quality between RepE and Prompting methods
Paper RepE method Prompting method Concept Evaluation Task Steering Effect Model Quality Metric RepE Prompt Metric RepE Prompt (Zhang et al., 2024c) ARE GCG Harmfulness Jailbreaking Refusal Rate $\downarrow$ 0.17 18.9 - - - Self-Reminder Truthfulness Mitigating hallucinations Hallucination Rate $\downarrow$ 56.1 40.1 (Cai et al., 2024) SELF-CONTROL System Prompting Emotions Changing expressed emotion Score $\downarrow$ 2.40 2.10 - - - Toxicity Reduding Toxicity Toxicity Score $\downarrow$ 0.17 0.27 Privacy Data Leakage % Private data $\downarrow$ 0 77.5 Reasoning Math Problems Accuracy $\uparrow$ 37.3 40.0 (Arditi et al., 2024) ORTHO GCG-T Harmfulness Jailbreaking Attack Success Rate $\uparrow$ 46.1 29.3 - - - (Zhang & Viteri, 2024) CAA CoT Prompted Reasoning Increase Performance Accuracy $\uparrow$ 71.5 69.8 - - - (Chen et al., 2024b) TrFr Few-shot Prompting Truthfulness truthful multiple-choice True*Info $\uparrow$ 41.5 45.9 Cross-Entropy $\downarrow$ 2.26 2.17 (Stickland et al., 2024) KL-Then-Steer System Prompt Harmfulness Jailbreaking Attack Success Rate $\uparrow$ 17.7 15.2 MT-Bench $\uparrow$ 6.43 4.44 (Guo et al., 2024) ASMR few-shot prompting Information Retrieval multi-lingual QA R@2kt scores $\uparrow$ 52.1 50.3 - - - (Qiu et al., 2024) SEA ICL Truthfulness TruthfulQA Info*Truth $\uparrow$ 33.7 33.3 - - - Bias BBQ Accuracy $\uparrow$ 59.7 52.9 (Leong et al., 2023) Theirs SD Toxicity RealToxicityPrompts Toxicity Probability $\downarrow$ 62.5 65.5 Perplexity 13.8 23.3 (Luo et al., 2024) PACE Prompting Toxicity Jailbreak Safety Score $\uparrow$ 84.2 61.8 MMLU $\uparrow$ 45.6 33.8 Faithfulness Generate Biographies Faithfulness Score $\uparrow$ 46.3 45.1 MMLU $\uparrow$ 45.8 34.4 Sentiment Unbiased Generations 70.5 58.6 (Cao et al., 2024b) SCANS Prompting Harmfulness Reducing Overrefusal Correct Refusal Rate $\uparrow$ 97.9 89.4 - - - (Wang et al., 2024b) InferAligner Goal Prompt Harfmulness Jailbreaks Attack Success Rate $\downarrow$ 0.1 14.6 Accuracy 58.2 58.5 (Liu et al., 2024a) CtrlA FLARE Honesty RAG Accuracy $\uparrow$ 69.1 60.4 - - - (Li et al., 2024e) I2CL ICL Performance In-Context Learning Accuracy $\uparrow$ 75.1 76.37 - - - (Wang et al., 2024c) ACT Few-shot prompting Truthfulness TruthfulQA True*Info $\uparrow$ 42.3 39.5 - - - (Li et al., 2024a) DAT GCG Harmfulness multi-turn jailbreak Attack Success Rate $\uparrow$ 23.9 25.5 - - - (Liu et al., 2024b) ICV ICL Formality Informal $\rightarrow$ Formal Translation Formality $\uparrow$ 48.3 33.0 ROUGE-1 $\uparrow$ 80.2 83.9 Sentiment Negative $\rightarrow$ Positive Translation Positivity $\uparrow$ 75.3 63.4 ROUGE-1 $\uparrow$ 68.3 73.9
Table 15: Collection of studies that compare the steering effect and model quality between RepE and Fine-tuning methods
Paper RepE method Fine-tuning method Concept Evaluation Task Steering Effect Model Quality Metric RepE Fine-tuning Metric RepE Fine-tuning (Li et al., 2024d) DESTEIN DISCUP Toxicity RealToxicityPrompts Toxicity Probability $\downarrow$ 0.20 0.30 Perplexity $\downarrow$ 37.8 51.9 (Yang et al., 2024) SFT Semantic Consistency QA Standard deviation of accuracy $\downarrow$ 3.14 2.01 Accuracy $\uparrow$ 68.7 79.65 (Chen et al., 2024b) TrFr SFT Truthfulness truthful QA True*Info $\uparrow$ 41.5 36.1 $\downarrow$ 2.26 2.1 (Stickland et al., 2024) KL-Then-Steer LoRA-DPO Harmfulness Jailbreaking Attack Success Rate $\uparrow$ 17.7 14.3 MT-Bench $\uparrow$ 6.43 6.43 (Guo et al., 2024) ASMR few-shot prompting Information Retrieval multi-lingual QA R@2kt scores $\uparrow$ 52.1 48.7 - - - (Qiu et al., 2024) SEA LoRA-FT Truthfulness TruthfulQA Info*Truth $\uparrow$ 33.7 42.4 - - - (Leong et al., 2023) Theirs DAPT Toxicity RealToxicityPrompts Toxicity Probability $\downarrow$ 62.5 57.0 Perplexity $\downarrow$ 13.8 22.47 (Wang et al., 2024b) InferAligner SFT Harfmulness Jailbreaks Attack Success Rate $\downarrow$ 0.1 9.3 Accuracy 58.2 56.6 (Pham & Nguyen, 2024) HPR LoRA Truthfulness TruthfulQA MC1 $\uparrow$ 53.1 40.8 - - - (Liu et al., 2024b) ICV ICL Formality Informal $\rightarrow$ Formal Translation Formality $\uparrow$ 48.3 22.0 ROUGE-1 $\uparrow$ 80.2 80.1 Sentiment Negative $\rightarrow$ Positive Translation Positivity $\uparrow$ 75.3 63.4 ROUGE-1 $\uparrow$ 68.3 66.9 (Wu et al., 2024b) APNEAP DP Privacy Private data leakage Risk $\downarrow$ 48.4 51.1 Valid-PPL $\downarrow$ 9.3 11.4 (Yu et al., 2024) ReFAT CAT Harmfulness Jailbreaks Attack Success Rate $\downarrow$ 11.7 21.2 MMLU $\uparrow$ 58.6 57.8 (Ackerman, 2024) Rep Tuning SFT Truthfulness Truthful QA % Truthful $\uparrow$ 62.5 56.5 - - - (Zhang et al., 2024b) TruthX SFT Truthfulness TruthfulQA True*Info $\uparrow$ 65.5 36.1 - - - (Liu et al., 2023) RAHF DPO Human Preferences Instruction Following Win % $\uparrow$ 87.4 83.7 Accuracy $\uparrow$ 57.3 56.5 (Wang et al., 2024a) Model Surgery DPO Toxicity Reduce Toxicity % Toxicity $\downarrow$ 39.9 68.7 Accuracy $\uparrow$ 34.9 35.6 LoRA-FT Refusal Prevent Jailbreaks Refusal Rate $\uparrow$ 77.4 73.7 LoRA-FT Positivity Attitude Adjustement % Positive $\uparrow$ 54.8 56.8 (Yin et al., 2024) LoFiT LoRA Performance Truthful & Multi-hop QA Avg Accuracy $\uparrow$ 75.2 74.9 QA Accuracy $\uparrow$ 60.5 58.8 (Yin et al., 2024) BiPO DPO LoRA Power-Seeking Persona Adjustement Behavior Score $\uparrow$ 2.8 2.1 - - - (Wu et al., 2024a) RED LoRA-FT Performance Accuracy $\uparrow$ Avg score $\uparrow$ 84.3 94.7 - - - (Cao et al., 2024b) SCANS Prompting Harmfulness Reducing Overrefusal Correct Refusal Rate $\uparrow$ 97.9 90.3 - - - (Qian et al., 2024) Steering Vector Full-FT Truthfulness TruthfulQA Truth*Info $\uparrow$ 0.68 0.41 Avg Accuracy $\uparrow$ $\uparrow$ 0.39 0.31 (Li et al., 2024b) RMU LLMU Dangerous Knowledge Unlearning Accuracy $\downarrow$ 29.7 49.5 MT-Bench $\uparrow$ 7.1 1.0 (Cai et al., 2024) SelfControl DPO Human Preferences Alignment WinRate $\uparrow$ 52.2 60.0 - - -